# Folded Context Condensation in Path Integral Formalism for Infinite Context Transformers

### v3

@@ I Introduction
Another critical limitation of Transformers is their computational inefficiency when handling long sequences. The self-attention mechanism inherently scales quadratically with sequence length in both memory and computation, making it impractical to process extensive contexts within a single pass. This limitation hinders the retention of dependencies across long documents, dialogues, or continuous interactions, prompting the development of alternative strategies such as Infini-Attention [7], which restructures the self-attention mechanism by introducing a hybrid memory system that allows interaction between local context tokens and external memory components, leveraging a linear attention [8] framework to reduce computational overhead. Another innovative solution is Titans [9], which incorporates trainable neural long-term memory, enabling the model to retain and retrieve past contextual information beyond traditional Transformer constraints. However, these methods often introduce external memory components that do not engage with the Feed-Forward Network, which we believe may reduce their ability to convey context information across segments.
@@
Building on this theoretical foundation in Section II, we present an implementation in Section III that extends Transformer architectures with a time-evolving memory buffer. This structured approach ensures that past token representations contribute constructively to future states, akin to propagators in quantum field theory. Our proposed model enables efficient long-range dependency modeling while significantly reducing memory consumption compared to conventional Transformers.
In section IV, we validate our approach through empirical evaluations, including the Passkey retrieval and summarization tasks. Our experimental results demonstrate that the proposed method achieves performance on par with the base model in contextual recall, while delivering superior memory efficiency compared to the base Llama-3.2 models [13].
@@ II Generalization of Transformer in Path Integral framework
$X_{i}^{\prime}$ $={\rm{softmax}}\left[\frac{X_{i}W_{Q}W_{K}^{T}}{\sqrt{d_{E}}}\begin{bmatrix}\cdots&X^{T}_{k-1}&X^{T}_{k}&\cdots\end{bmatrix}\right]$
@@
where $FF$ stands for Feed-Forward Network and $W_{Q,\,K,\,V}\in\mathbb{R}^{\,d_{E}\times d_{M}}$ are query, key and value matrices of the Transformer. Here we ignored multi-head attention for simplicity.
We introduced the concept of time into the Transformer structure by generalizing Transformers in the Path Integral framework and finding correspondences between them. The Transformer can be reinterpreted as the temporal evolution of an input token, and attention can be understood as a transition of an input token state to all possible token states in the same timeline. As the input token passes through each layer of Transformers, the time evolution is achieved, which can be interpreted as traversing all possible positional and temporal paths. Then, by reducing the cross entropy loss between targets and the output of Transformers, Transformers are trained to find the Least Action path of the Path Integral, which can be captured by the following equation as
@@
which cross-entropy loss of the Transformers is calculated by the transition amplitude between two states through path-integral. As the model is trained, the cross-entropy loss becomes decreasing. Then, to get larger prediction probability as smaller cross-entropy loss, the paths with high frequency are canceled out so that the only paths carrying low valued-actions survive, which achieves the Least Action Principle. This may explain why attention-based Transformers perform so well when estimating the next token state. In the next section, we will introduce an application from the perspective of time evolution.
III Folded context Condensation: the implementation of the model
@@
We consider a certain case of combining outputs of Transformer layers for separated input segments. A length $L$ sequence is divided into segments of length $l$, such that $N_{s}=L/l$, and then the tokens in each segment are put into the Transformer layers sequentially. Here, we suggest how to deliver the information of a segment to another segment in the Transformer architecture. In the previous section, by reinterpreting the Transformers in the Path Integral framework, we find that the Feed-Forward Network makes the temporal evolution of the token state. Therefore, it is expected that there can be a phase difference between the token states in separated segments.
As shown in Fig. 2, we define a token state $|x_{i}^{s},\,t_{n}^{s}\rangle$ on the position $i$ in $s$-th segment and the token state evolves in a Transformer layer as
@@
${\cal N}\sum_{k}|x_{k}^{s},\,t_{n}^{s}\rangle\langle x_{k}^{s},\,t_{n}^{s}|x_{i}^{s},\,t_{n}^{s}\rangle$ (13)
with $i=1,\,\cdots,\,l$ and $s=1,\,\cdots,\,N_{s}$. The temporal evolution of the token state makes the phase shift of the token state, and each layer gives a different phase shift. If the phase difference between different segments is not properly handled, tokens from separated segments may interfere destructively, leading to the loss of important historical information from previous segments. Here is a solution to make the constructive interference between the segments, that is we impose a time gap between the segments to be periodical, which can be accomplished by applying recurrent Transformer layer with the segments. For example, we input $s-1$-th and $s$-th segments in parallel into a Transformer layer as
@@
$\qquad\left.+\left|x_{k}^{s},\,t_{n}^{s}\rangle\langle x_{k}^{s},\,t_{n}^{s}\right|\right)|x_{i}^{s},\,t_{n}^{s}\rangle\,,$ (14)
where we apply the boundary condition as
$t_{n}^{s}=t_{n}^{s-1}+\Delta t$ (15)
by concatenating the output of the layer for $x^{s-1}$ with the input of the layer for $x^{s}$ step in Fig. 2. In this way, every segment will line up with an integer-multiple time gap so that every segment contributes to constructive interference.
This periodic temporal evolution and the transition between $s-1$-th and $s$-th segments can be presented as
$X^{s\,\prime}=\rm{softmax}\left(\frac{Q^{s}}{\sqrt{d_{E}}}\begin{bmatrix}K^{s-1\,T}&K^{s\,T}\end{bmatrix}\right)\left(\begin{matrix}V^{s-1}\\ V^{s}\end{matrix}\right)W_{FF}$ (16)
in a Transformer layer shown as Fig. 2, where $Y^{s}=X^{s}W_{Y}$ with $X^{s}\in\mathbb{R}^{\,l\times d_{E}}$ and $Y=Q,\,K,\,V$. The masking for the attention matrix is given in Fig. 3. We note that the all dot-product attention between $Q^{s}$ and $K^{s-1}$ are not masked to relate all contextual information from past segments to the current query. Then, every layer has a recurrent structure transferring the information from $s-1$-th to $s$-th segments.
@@
We explain how we can implement the model to reduce resources and apply it to multi-layer Transformers. When training Transformers, the attention mechanism is one of the most resource-consuming components, scaling quadratically with the input length. So, we focus on how we reduce the token length for attention while preserving the quality of the prediction by conserving the historical information through the sequence divided by the segments. We suppose that the token sequence is divided by $N_{s}(L/l)$ segments having the same token length $l$ as in the previous section, and the segments are sequentially put into the multi-layer Transformers as shown in Fig. 4. First, token embedding is performed for each segment, where each has a different position ids. Then, the $s$-th segment is concatenated with the previous recurrent $s-1$-th segment to be put into the attention as length $2l$ key and value vectors while the query has length $l$ as in Fig. 2. After the output of the attention is added up with the residual unit, it is put into the Feed-Forward Network. Finally, the outputs of the multi-layer Transformers are matched with each target for each corresponding segment to get the cross-entropy loss which is summed up as
@@
where $L_{s}$ is a cross-entropy loss for $s$-th segment. The model is trained to reduce the total cross-entropy loss, $L_{\rm{total}}$.
@@ IV Experiments
IV-A Memory usage measurement & Speed Comparison
@@
IV-B Passkey retrieval task

### v4

@@ I Introduction
Another critical limitation of Transformers is their computational inefficiency when handling long sequences. The self-attention mechanism inherently scales quadratically with sequence length in both memory and computation, making it impractical to process extensive contexts within a single pass. This limitation hinders the retention of dependencies across long documents, dialogues, or continuous interactions, prompting the development of alternative strategies such as Infini-Attention [7], which restructures the self-attention mechanism by introducing a hybrid memory system that allows interaction between local context tokens and external memory components, leveraging a linear attention [8] framework to reduce computational overhead. Another innovative solution is Titans [9], which incorporates trainable neural long-term memory, enabling the model to retain and retrieve past contextual information beyond traditional Transformer constraints. However, these methods often introduce external memory components that do not engage with the Feed-Forward Network, which we believe it may reduce their ability to convey context information across segments.
@@
Building on this theoretical foundation in Section II, we present an implementation in Section III and IV that extends Transformer architectures with a time-evolving memory buffer. This structured approach ensures that past token representations contribute constructively to future states, akin to propagators in quantum field theory. Our proposed model enables efficient long-range dependency modeling while significantly reducing memory consumption compared to conventional Transformers.
In section V, we validate our approach through empirical evaluations, including the Passkey retrieval and summarization tasks. Our experimental results demonstrate that the proposed method achieves performance on par with the base model in contextual recall, while delivering superior memory efficiency compared to the base Llama-3.2 models [13].
@@ II Generalization of Transformer in Path Integral framework
$X_{i}^{\prime}$ $={\rm{softmax}}\left[\frac{X_{i}W_{Q}W_{K}^{T}}{\sqrt{d_{M}}}\begin{bmatrix}\cdots&X^{T}_{k-1}&X^{T}_{k}&\cdots\end{bmatrix}\right]$
@@
where $FF$ stands for Feed-Forward Network and $W_{Q,\,K,\,V}\in\mathbb{R}^{\,d_{E}\times d_{M}}$ are query, key and value matrices of the Transformer. That is the generalized form of decoder-only Transformer. Here we ignored multi-head attention for simplicity.
We introduced the concept of time into the Transformer structure by generalizing it within the Path Integral framework and establishing formal correspondences. This reinterpretation allows us to view the Transformer mechanism as the temporal evolution of an input token state. In this perspective, attention serves as a transition mechanism, enabling an input token to explore all possible token states within the same timeline. As the input token propagates through successive Transformer layers, its evolution follows a trajectory that encompasses all potential positional and temporal paths.
Training the Transformer can then be understood as an optimization process that identifies the Least Action path within this Path Integral formulation. Specifically, minimizing the cross-entropy loss between the predicted outputs and target tokens can be interpreted as selecting the optimal trajectory that satisfies the Principle of Least Action. Mathematically, this is expressed as
@@
Here, the transition amplitude $\mathcal{M}_{ij}$ represents the probability amplitude of moving from the initial state $|j\rangle$ to the final state $|i\rangle$, summing over all possible evolutionary pathways governed by the Lagrangian $\mathcal{L}$. The Transformer training process thus naturally suppresses high-frequency paths that contribute minimally to the model’s final output while reinforcing trajectories that minimize the action. This mechanism leads to the emergence of the Least Action Principle in Transformer optimization:
$-\ln\left|\mathcal{M}_{ij}\right|^{2}\rightarrow 0\,.$ (13)
Consequently, this framework provides an intuitive explanation for the superior performance of attention-based Transformers in next-token prediction. By dynamically refining token transitions through a learned path integral landscape, the Transformer effectively selects the most efficient computational paths that yield coherent and contextually relevant sequences.
In the next section, we will explore practical applications of this perspective in modeling time evolution within neural architectures.
III Folded context Condensation
@@
In our Transformer architectures, token states undergo temporal evolution across layers, leading to phase shifts that influence coherence in long-range dependencies. If unregulated, these shifts may induce destructive interference between token states of separate segments, causing information loss. To address this, we analyze the impact of token state phase shifts and introduce a periodic phase gap mechanism that ensures constructive interference.
As depicted in Fig. 2, the token state at position $i$ in segment $s$ evolves temporally through the Transformer layer as
@@
$\quad{\cal N}\sum_{k}|x_{k}^{s},t_{n}^{s}\rangle\langle x_{k}^{s},t_{n}^{s}|x_{i}^{s},t_{n}^{s}\rangle\,.$ (14)
This equation suggests that each Transformer layer, after training, exhibits distinct phase evolution profiles aligning with eigenmodes of token state transitions. The layer parameters adapt to minimize phase misalignment, leading to the emergence of layer-specific eigen-frequencies.
This phenomenon is analogous to resonant systems such as stringed or wind instruments, where standing wave patterns form due to imposed boundary conditions. Consequently, the phase evolution of token states in a Transformer layer follows the wave equation:
$i\hbar\frac{\partial}{\partial t}\Psi_{i}^{s}-\mathcal{H}\Psi_{i}^{s}=0,$ (15)
where $\Psi_{i}^{s}$ denotes the token state at position $i$ in segment $s$, and $\mathcal{H}$ represents the Hamiltonian of the layer’s transformation. This equation leads to standing wave solutions constrained by the boundary conditions between input and output tokens, indicating that token states evolve according to intrinsic phase shifts governed by learned parameters. These parameters are trained to align phase shifts with resonant frequencies, $f_{k}$’s, at each layer:
$\Delta\phi_{n}=\omega_{n}(\cdots,\,f_{k},\cdots)\Delta t_{n},\quad\omega_{n}\sim F(\mathcal{H},\,\Psi_{i}^{s}),$ (16)
where $\omega_{n}$ is the overall frequency, a function of resonant frequencies, and given by the input token state $\Psi_{i}^{s}$ applied by Hamiltonian. We note that $\Delta t_{n}$ is the time period of the standing wave at the $n$-th layer. After training, these values are fixed to ensure that token state transitions reinforce the context established by prior layers rather than interfere destructively.
To prevent destructive interference, we introduce a periodic phase gap between segments, enforcing coherence in token state transitions. Misaligned phase differences between adjacent segments can introduce interference effects, hampering information retention. We define the segment-to-segment transition as
@@
$\qquad\left.+\left|x_{k}^{s},t_{n}^{s}\rangle\langle x_{k}^{s},t_{n}^{s}\right|\right)|x_{i}^{s},t_{n}^{s}\rangle.$ (17)
By explicitly incorporating prior segment states into the transition, the alignment of segments at integer-multiple time gaps,
$t_{n}^{s}=t_{n}^{s-1}+\Delta t_{n},$ (18)
enables constructive interference, preserving historical context across segments.
To implement this in practice, we propose a recurrent segment processing approach wherein token states from previous segments are periodically reinjected into the Transformer as shown in Fig. 2. This structured recurrence ensures phase-aligned reinforcement, analogous to standing wave resonances. By enforcing these phase-aligned boundary conditions, the architecture sustains long-range dependencies through constructive interference, mitigating contextual degradation over extended sequences.
IV The implementation of the model
@@
Building upon the structured recurrence and phase coherence mechanisms introduced in the previous sections, we now formalize the implementation of segmental processing in Transformer architectures. Our approach optimizes memory efficiency while preserving long-range dependencies by leveraging recurrent attention across segment transitions.
To ensure seamless information propagation across segments, we define the periodic temporal evolution of token states in the Transformer layer as:
$X^{s\,\prime}=\rm{softmax}\left(\frac{Q^{s}}{\sqrt{d_{M}}}\begin{bmatrix}K^{s-1\,T}&K^{s\,T}\end{bmatrix}\right)\left(\begin{matrix}V^{s-1}\\ V^{s}\end{matrix}\right)W_{FF},$ (19)
where $Y^{s}=X^{s}W_{Y}$ with $X^{s}\in\mathbb{R}^{l\times d_{E}}$ and $Y=Q,\,K,\,V$. The attention mask structure, depicted in Fig. 3, ensures that the query $Q^{s}$ retains full access to past segment keys $K^{s-1}$ while maintaining causal restrictions within the current segment $K^{s}$. By enabling unmasked attention between $Q^{s}$ and $K^{s-1}$, we ensure that past contextual information remains accessible, reinforcing coherence between segments.
Given that self-attention mechanisms scale quadratically with input length, reducing computational overhead is crucial for efficient training and inference. Our recurrent segment-wise approach mitigates the resource constraints associated with large-scale Transformers while ensuring robust sequence modeling. We achieve this by partitioning an input sequence of length $L$ into $N_{s}$(= $L/l$) segments, each of uniform length $l$, and sequentially processing them through the multi-layer Transformer as shown in Fig. 4.
During training, token embeddings are first assigned distinct positional indices per segment. The $s$-th segment is then concatenated with the recurrent representation from the $(s-1)$-th segment, forming a combined key-value set of length $2l$, while the query remains restricted to $l$ positions, as described in Fig. 2. This formulation effectively captures and propagates historical information while limiting redundant computation.
Following self-attention, the output undergoes residual connection processing and is passed through a Feed-Forward Network (FFN). The final layer-wise outputs align with corresponding segment-wise targets, yielding the total loss function:
@@
where $L_{s}$ represents the cross-entropy loss for the $s$-th segment. The model is optimized to minimize $L_{\rm{total}}$, ensuring robust long-range dependency retention while maintaining computational efficiency.
@@ V Experiments
V-A Memory usage measurement & Speed Comparison
@@
V-B Passkey retrieval task